A diffraction grating has 3 200 rulings/cm. On a screen 2.50 m from the grating, it is found that for a particular order m, the maxima corresponding to two closely spaced wavelengths of sodium (589.0 nm and 589.6 nm) are separated by 1.21 mm. Determine the value of m.
m = ____

m is not 3,4 or 5

PLEASE ONLY ANSWER IF YOU KNOW ITS CORRECT
Answer #1

Diffraction grating, d = 3200 rulings/cm = [(0.01) / (3200)] m

d = 3.12 x 10-6 m

Under the assumption that sin \theta\approx tan \theta = (y / L)

We know that, \lambda1 correspond to the wavelength that is higher on the screen, while \lambda2 is lower on the screen.

For a bright fringes, we have

d sin \theta1 = m \lambda1

d (y1 / L) = m \lambda1                                                  { eq.1 }

And

d sin \theta2 = m \lambda2

d (y2 / L) = m \lambda2

d [y1 - (1.21 x 10-3 m)] / L = m \lambda2                                           

d (y1 / L) - (1.21 x 10-3 m) (d / L) = m \lambda2                                              { eq.2 }

inserting the value of eq.1 in eq.2 & we get -

m \lambda1 - (1.21 x 10-3 m) (d / L) = m \lambda2   

m (\lambda1 - \lambda2) = (1.21 x 10-3 m) (d / L)

m [(589.6 - 589) x 10-9 m] = [(1.21 x 10-3 m) (3.12 x 10-6 m)] / (2.5 m)

m = (1.51008 x 10-9 m) / (0.6 x 10-9 m)

m = 2.5